OpenStudy (anonymous):

Arianna kicks a soccer ball off the ground and into the air with an initial velocity of 42 feet per second. Assume the starting height of the ball is 0 feet. Approximately, how long does it take until the soccer ball hits the ground again? 1.3 secs 2.6 secs 16.0 secs 27.6 secs

OpenStudy (amistre64):

what do we know that will helps us with this?

OpenStudy (amistre64):

and is the initial velocity vertical? or angled?

OpenStudy (anonymous):

vertical i tink

OpenStudy (amistre64):

i think so too, its hard to come to a definitive result otherwise

OpenStudy (amistre64):

this tends to be more of a physics question, seeing as how if it was mathical they provide more of a formula to solve

OpenStudy (amistre64):

since our velocity is in feet; lets use gravity as -32 ft/sec/sec

OpenStudy (amistre64):

this is a constant acceleration due to gravity on earth: a(t) = -32 to get to velocity we need to "antiderive" or inegrate it up. \[\int -32\ dt=-32t+C\]

OpenStudy (amistre64):

that was odd ... computer went a little whacky

OpenStudy (amistre64):

at time = 0; or t=0 we have a velocity of 42 v(0) = -32(0) + C = 42 C = 42 v(t) = -32t +42 .... so far so good

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

to get to position, or height, we need to up this one more time

OpenStudy (amistre64):

\[\int -32t+42\ dt=-16t^2+42t+C\] and since at t=0 we have a height of 0 ... h(0) = -16(0)^2 +42(0) + C = 0 C = 0 h(t) = -16t^2 +42t now we can solve for "t" that will give us a height of 0 ... you know, other than at t=0 :)

OpenStudy (amistre64):

h(t) = -16t^2 +42t = 0 2t(-8t +21) = 0 t = 0 or t = 21/8, 2 5/8

OpenStudy (amistre64):

id say 2.6 is a good bet

OpenStudy (anonymous):

thank u very much

OpenStudy (amistre64):

youre welcome :)